1.    Mike, Rosie, and Shobber live in separate houses along a dark and windy road. The following represent their marginal benefits for street lights:

MBMike=200-2QM

MBRosie=100-QR

MBShobber=100-2QS

where QM represents the quantity of street lights consumed by Mike, QR is the quantity of street lights consumed by Rosie and QS is the quantity of street lights consumed by Shobber. The Mayor of their town considers street lights to be a public good and is charged with purchasing the optimal number of street lights from Boone’s Light Shop. Boone’s is willing to sell street lights for $150 per light.

a.     In what sense, if any, do street lights qualify as a public good? What is the relationship between QM, QR, and QS (i.e., greater than, less than or equal to) likely to be when the Mayor installs the lights? Why?





1
Expert's answer
2021-10-18T09:07:21-0400

"MB_{Mike}=200-2Q_M \\

MB_{Rosie}=100-Q_R \\ 

MB_{Shobber}=100-2Q_S"

street lights = $150

Total marginal benefits = "MB_{Mike} + MB_{Rosie} + MB_{Shobber}"

"= 200-2Q_M + 100- Q_R + 100 -2Q_S \\

= 400 -5Q"